Can you help me please to simplify these two negative exponents.
8 - 2

4 - 2

Found 2 solutions by richwmiller, rothauserc:
Answer by richwmiller(17219)

If you mean
8^-2
and
4^-2
8^-2=1/8^2=1/64
4^-2=1/4^2=1/16

Answer by rothauserc(4718)

8^-2 = 1 / 8^2 = 1 / 64
4^-2 = 1 / 4^2 = 1 / 16